# PROSPECT: Precision Robot Spectroscopy Exploration and Characterization Tool

### v1

@@ Abstract
Near Infrared (NIR) spectroscopy is widely used in industrial quality control and automation to test the purity and material quality of items. In this research, we propose a novel sensorized end effector and acquisition strategy to capture spectral signatures from objects and register them with a 3D point cloud. Our methodology first takes a 3D scan of an object generated by a time-of-flight depth camera and decomposes the object into a series of planned viewpoints covering the surface. We generate motion plans for a robot manipulator and end-effector to visit these viewpoints while maintaining a fixed distance and surface normal to ensure maximal spectral signal quality enabled by the spherical motion of the end-effector. By continuously acquiring surface reflectance values as the end-effector scans the target object, the autonomous system develops a four-dimensional model of the target object: position in an $R^{3}$ coordinate frame, and a wavelength vector denoting the associated spectral signature. We demonstrate this system in building spectral-spatial object profiles of increasingly complex geometries. As a point of comparison, we show our proposed system and spectral acquisition planning yields more consistent signal signals than naïve point scanning strategies for capturing spectral information over complex surface geometries. Our work represents a significant step towards high-resolution spectral-spatial sensor fusion for automated quality assessment.
@@ I Introduction
Quality, repeatability, and safety are key considerations for manufacturing processes. In a zero-defect manufacturing paradigm, items are selected for grading and evaluation at multiple points in the production process, with a preference for non-destructive testing when possible [1]. Spectroscopy is a commonly used non-contact technique to assess manufacturing quality. By analyzing patterns of absorption or emission of electromagnetic radiation by a material, spectroscopy can provide information about its composition, structure, and physical properties [2]. Spectral information can alternatively be acquired through hyperspectral imaging (HSI) – a technique that extends RGB imaging to $n$ discrete wavelengths. Spectral sensing has many diverse applications in agriculture [3], food processing [4], geology [5], and remote sensing [6].
Spectral data by itself is useful in determining the material nature of items with precision far exceeding RGB imaging, and without physical contact. Nonetheless, fusing spectral data with spatial information creates a richer understanding of objects. For example, knowing where agricultural items are bruised or diseased can allow these regions to be excised, salvaging the remaining product. Additionally knowing where specific corrosion patterns exist on a built structure can yield an estimation of the extent of structural decay.
@@
However, manual acquisition of spectral information is time intensive and requires precision alignment of a probing element to cover a surface. Similarly, HSI is constrained in its ability to only collect data in a 2D projection. The majority of HSI systems use a pushbroom acquisition, where the camera or object must be linearly translated to acquire a full image. Additionally, HSI systems require intense active illumination for indoor operation. While point spectroscopy only acquires a 1D spectral profile from the target object, it can offer an order of magnitude increase in spectral resolution and cost-effectiveness at these points.
Acquiring multiple spectral measurements from over a surface requires a probe to be precisely translated, as signal formation is highly dependent on the distance of the sensed object from the measurement device. Thus the core question of our work is: Can point spectroscopy coupled with a dexterous end-effector autonomously generate high resolution spectral-spatial profiles of objects for inspection?
@@
  1. 1.
    Design, kinematics modeling, and validation of an end-effector for precise spectral measurements.
  2. 2.
    An algorithm for spectral viewpoint selection to maximize signal consistency and coverage over a 3D surface
  3. 3.
@@ II-A Spectral Quality Inspection
Spectral sensing has enjoyed widespread use in many industries, but chiefly in the realm of precision agriculture. The presence of the red-edge effect, caused by chlorophyll reflectance between 690-730 nm, has long been used to quantify the health of vegetation [7]. By analyzing various parts of the reflectance spectrum, researchers developed methods to detect plant ripeness [8], stress [9], disease [10], phenotype [11], and invasive species [12].
Classification of spectral signals via machine learning has been demonstrated in a variety of use cases including pharmaceutical [13], medical [14], and recycling [15]. The food industry relies on NIR spectroscopy to estimate the abundance of sugar, fat, and overall quality of foods [16, 17, 18]. In these approaches, spectral signatures from a point spectrometer or imaging array are passed into a model that generates class probabilities representing the likelihood of the current signal belonging to each class. Another machine learning strategy attempts to compare unknown spectral signatures against a dictionary of reference spectra. These libraries have been extensively compiled with thousands of known materials and objects and are independent of the measuring spectrometer [19].
@@ II-B Spectral-Spatial Point Clouds
Spectral information alone is not necessarily enough to adequately describe a complex object. HSI datacubes contain 2D spatial information, but constrain objects to a single plane of acquisition. Combining spectral information with point clouds allows a geometric interpretation of material information. [20] showed significant improvements in terrain classification with remotely sensed hyperspectral and LIDAR measurements when fusing the two modalities. [21, 22, 23] fused snapshot hyperspectral data with LIDAR to create spatial-spectral digital surface models, but only with objects at a distance. [24] also created hyperspectral point cloud data, but used an arm-mounted HSI to acquire hyperspectral images with optimizations for successive arm placements [25]. To the best of our knowledge, no general approach has been developed to fuse high-resolution point-spectral measurements with points clouds at close distances.
@@ II-C Robot Spectroscopy
The use of spectroscopy in robotics has gained traction as a way to classify household objects [26], survey geologic sites [27], explore underwater [28], and assess fruit ripeness [29]. Our previous work has been concerned with making pre-grasp inferences using single point spectral measurements [30, 31, 32]. However, in all these works we have made assumptions of a fixed operating distance from object to sensor aperture, which simplified our modeling of spectral signals. The design and operation of these sensors assumed a single point scan was representative of the entire item composition, which is true for select tasks in object manipulation. For understanding the spectral signatures of complex objects in automation and grading processes, we can no longer assume uniformity.
@@ III System Design
PROSPECT is a robotic spectroscopy system that incorporates mechanical and optical elements for accurate formation and registration of spectral signatures with scanned geometries. PROSPECT uses a miniaturized 6-degree-of-freedom (DoF) Stewart platform for precision orientation of a spectroscopy sensor suite relative to a measured surface. In this study, a 6-DoF UR3e robotic arm provides global motion around an object, while PROSPECT provides fine surface normalization and offsetting.
@@ III-A Mechanical
PROSPECT is designed to function as a precision end-effector adaptable to global motion platforms such as robot arms, gantries, or drones. PROSPECT contains a sensor suite with a fiber-optic cable, illuminating lights, and time-of-flight-sensors. These are mounted on the top platform of the Stewart mechanism, which allows both spherical motion (for surface normal matching) and linear motion (for surface offsetting).This enables fine adjustment of the fiber-optic cable without needing to move the full weight of the global motion stage. The smoothness of spherical rotation also minimizes damage to the fiber optic cable which cannot be pinched or severely curved without permanent damage.
Figure 2: (Left) PROSPECT end-effector component diagram showing subcomponents. (Right) Identification of parameters and axes for kinematics and motion planning.
The PROSPECT end-effector is composed of a rigid base, a series of parallel links each composed of a horn $l_{h}$ and an arm $l_{a}$, and a sensorized platform as shown in Fig. 2. The base is 3D printed in Onyx on a Markforged X7 and contains wire routing and mounting to connect to a standard interface (Universal Robotics). Six XL330-M288-T servo motors (Dnyamixel) are mounted to the base. Each servo is oriented radially outward and positioned at alternating small and large angle increments to maximize the motor packing density. The small angle between neighboring servos, $\theta_{b}$, is measured from the $z$ axis to the virtual point at the intersection between the servo axis and the plane of the arms.
The base also holds a small 3D printed electronics enclosure with a microcontroller (Adafruit QtPy), a servo controller, and servo power board. Attached to each motor is a custom servo horn leading to the anchor point of ball-ended arm. Each ball end arm consists of one left and one right threaded rod end ball joint joined by an aluminum turnbuckle to allow adjustment of the lengths of the arms for different desired workspaces. The upper ball joint of each arm connects to the outer edge of the top platform, as shown in Fig. 2 below. The angle between the platform connection points, $\theta_{p}$, is measured from the z axis to the center of the upper ball joint.
@@ III-B Optical
The top platform contains the sensors and electronics that enable point spectroscopy. Current spectrometers are too fragile and bulky to be fully encapsulated in the top platform; instead, the center of the platform contains a shielded fiber optic cable (ThorLabs) to guide light from the field of view of the platform and into a spectrometer at the base of the manipulator. The cable contains a 400 $\mu$m silica glass core with a numerical aperture (NA) of 0.50. The top platform clamps the ferrule end of the fiber to a fixed normal orientation relative to the platform surface. The fiber optic cable from the top plate connects to a VNIR diffraction grating spectrometer (Ibsen Photonics) with sensitivity from 500-1100 nm with 256 discrete wavelength samples. PROSPECT pairs with any grating spectrometer operating in the VIS-SWIR range (400-2200 nm); thus, it is reconfigurable for sensitivity to either a broad or narrow range of the electromagnetic spectrum given the dictates of the inspection application.
Before operation, the spectrometer setup is calibrated with a Spectralon reference standard and a dark noise collection according to [33].
@@
In eq. (1), the tilde operator refers to the median of the spectral samples, collected over 10 measurements. The subscript $t$ indicates all measurements were collected with the same integration period. ${\tilde{L}}_{spec,t}$ refers to the median measurement taken from the Spectralon target at the nominal viewpoint distance. ${\tilde{D}}_{spec,t}$ refers to the median measurement taken with the spectrometer capped to incoming light. This reading accounts for dark current and other thermally induced electrical noise. ${S_{spec,t}}$ refers to the raw spectral signal for a particular sample and ${S_{calspec,t}}$ is the calibrated output.
The top platform contains 3 miniaturized Quartz Tungsten Halogen lamps (Ocean Insight). The lamps bathe the target surface in full-spectrum illumination, providing sufficient signal for spectroscopic measurements. The lights operate at 5V and are controlled with a MOSFET circuit.
The top platform also contain 3 time of flight (ToF) distance sensors (Adafruit VL53L4CD) oriented at 120 degree increments offset from the QTH lamps. The ToF sensors emit a pulse with a half-angle of 9. Given their spacing, the platform may use the three sensors to approximate a surface normal to the target surface at distances less than 6.0 cm. Alternatively, these sensors may be used to find the distance from the the fiber ferrule tip to the object of interest.
@@ IV-A Inverse Kinematics
    Compute $\vec{q_{i}}$ from eq. (3) and $\vec{l_{i}}$ from eq. (4)
@@
    Solve for $\alpha_{i}$ using eq. (5)
@@ IV-B Validation of Inverse Kinematics
To fully quantify end-effector pose error, we report errors in the position and orientation. Position error is given as the root mean squared error (RMSE) between the motion captures observed 3D coordinate $(x_{1},y_{1},z_{1})_{k}$ and the commanded pose $(x_{2},y_{2},z_{2})_{k}$ for sample $k$ out of $n$ total randomly sampled poses. The RMSE metric is given as:
@@
Fig. 3 shows the results of the kinematics validation experiments. Overall, the system is found to have a kinematic accuracy of $3.73\pm 3.10$ mm in positional errors and $0.036\pm 0.025$ radians in rotational accuracy. The larger errors in pose arise when the system suffers from backlash as the system quickly rotates to a pose at the opposite of its feasible workspace. The system’s minimal angular error indicates precise surface normal matching is possible.
@@ V Viewpoint Planning
With a point spectrometer, only a small portion of the object of interest can be characterized in one scan. To characterize the entire object, we determine many points of interest covering the surface of the object that need to be viewed with the point spectrometer. We assume the existence of a point cloud which will form the basis for our viewpoint planning. To maximize scan consistency and spectral signal strength, the fiber optic cable should be placed at a fixed distance from points of interest, as well as normal to the surface at that point.
@@ V-A Point Cloud Preprocessing
The raw point cloud from the 3D ToF Sensor (Azure Kinect, Microsoft) mounted above the arm contains points belonging the object of interest, as well as other extraneous points. Open3d [36] is used for the point cloud processing. The extraneous points are removed through a series of preprocessing steps. First, the cloud is cropped based off of empirically determined global bounds. Then, the table upon which the object rests is discarded by removing inliers belonging to a plane found with RANSAC [37]. Finally, noisy points are removed, and points belonging to the largest object remaining in the scene are selected with Density Based Clustering [38].
@@ V-B Approach Positioning
Once the point cloud has been preprocessed, viewpoints are selected with Algorithm  1. Consistently spaced points that cover the viewable surface are found through Voxel based uniform down-sampling. The normal at each downsampled point is then estimated. For each point and corresponding normal, the Stewart platform is positioned at a fixed distance from the surface along the normal such that the fiber optic probe is pointed parallel to the normal, facing the point. The arm is positioned to a point with only a Z-axis offset to the Stewart platform, with a fixed rotation for all points, so that Algorithm  1 generalizes to the possible configuration of a Stewart platform affixed to the end of a 3-DoF system such as an $XYZ$ Gantry. Our experiments use a UR3e Cobot (Universal Robotics) with MoveIt motion planning backend to achieve the desired approach positions [39].

### v2

@@ Abstract
Near Infrared (NIR) spectroscopy is widely used in industrial quality control and automation to test the purity and grade of items. In this research, we propose a novel sensorized end effector and acquisition strategy to capture spectral signatures from objects and register them with a 3D point cloud. Our methodology first takes a 3D scan of an object generated by a time-of-flight depth camera and decomposes the object into a series of planned viewpoints covering the surface. We generate motion plans for a robot manipulator and end-effector to visit these viewpoints while maintaining a fixed distance and surface normal. This process is enabled by the spherical motion of the end-effector and ensures maximal spectral signal quality. By continuously acquiring surface reflectance values as the end-effector scans the target object, the autonomous system develops a four-dimensional model of the target object: position in an $R^{3}$ coordinate frame, and a reflectance vector denoting the associated spectral signature. We demonstrate this system in building spectral-spatial object profiles of increasingly complex geometries. We show the proposed system and spectral acquisition planning produce more consistent spectral signals than naïve point scanning strategies. Our work represents a significant step towards high-resolution spectral-spatial sensor fusion for automated quality assessment.
@@ I Introduction
Quality, repeatability, and safety are key considerations for manufacturing processes. In a zero-defect manufacturing paradigm, items are selected for grading and evaluation at multiple points in the production process, with a preference for non-destructive testing when possible [1]. Spectroscopy is a commonly used non-contact technique to assess manufacturing quality. By analyzing patterns of absorption or emission of electromagnetic radiation by a material, spectroscopy provides information about its composition, structure, and physical properties [2]. Spectral information is alternatively acquired through hyperspectral imaging (HSI) – a technique that extends RGB imaging to $n$ discrete wavelengths. Spectral sensing has many diverse applications in agriculture [3], food processing [4], geology [5], and remote sensing [6].
Spectral data are useful for determining the material nature of items with precision far exceeding RGB imaging, and without physical contact. Therefore, fusing spectral data with spatial information creates richer object models. For example, knowing where agricultural items are bruised or diseased allows these regions to be excised, preserving the remaining product. Additionally, knowing where specific corrosion patterns exist on a built structure helps to estimate the extent of structural decay.
@@
However, manual acquisition of spectral information is time intensive and requires precision alignment of a probing element to cover a surface. Similarly, HSI reduces spectral information to a 2D projection. Most HSI systems use a pushbroom acquisition, in which the camera or object must be linearly translated to acquire a full image. HSI systems also require intense active illumination for indoor operation. In contrast, point spectroscopy acquires a 1D spectral profile from the target object, but it can offer an order of magnitude increase in spectral resolution and cost-effectiveness at these points.
Acquiring multiple spectral measurements from over a surface requires a probe to be precisely translated as signal formation is highly dependent on the distance of the sensed object from the measurement device. Thus, the core question of our work is: Can point spectroscopy coupled with a dexterous end-effector autonomously generate high-resolution spectral-spatial profiles of objects for inspection?
@@
  • Design, kinematics modeling, and validation of an end-effector for precise point spectral measurements.
  • An algorithm for spectral viewpoint selection to maximize signal consistency and coverage over a 3D surface.
@@ II-A Spectral Quality Inspection
Spectral sensing has enjoyed widespread use in many industries, but chiefly in the realm of precision agriculture. The presence of the red-edge effect, caused by chlorophyll reflectance between 690-730 nm, has long been used to quantify the health of vegetation [7]. By analyzing various parts of the reflectance spectrum, researchers developed methods to detect plant ripeness [8], stress [9], disease [10], and phenotype [11].
Classification of spectral signals via machine learning has been demonstrated in a variety of use cases, including pharmaceutical [12], medical [13], and recycling [14]. The food industry relies on NIR spectroscopy to estimate the abundance of sugar, fat, and the overall quality of foodstuffs [15, 16, 17]. In these approaches, spectral signatures from a point spectrometer or imaging array are passed into a model that generates class probabilities representing the likelihood of the current signal belonging to each class. Another machine learning strategy attempts to compare unknown spectral signatures against a dictionary of reference spectra. These libraries have been extensively compiled with thousands of known materials and objects and are independent of the measuring spectrometer [18].
@@ II-B Spectral-Spatial Point Clouds
Spectral information alone is not necessarily enough to adequately describe a complex object. HSI datacubes contain 2D spatial information, but constrain objects to a single plane of acquisition. Combining spectral information with point clouds allows a geometric interpretation of material information. [19] showed significant improvements in terrain classification by fusing remotely sensed hyperspectral datacubes and with LIDAR point clouds. [20, 21, 22] fused snapshot hyperspectral data with LIDAR to create spatial-spectral digital surface models, but only with objects at a distance. [23, 24] also created hyperspectral point cloud data, but used an arm-mounted HSI to acquire hyperspectral images with optimizations for successive arm placements. To the best of our knowledge, no general approach has been developed to fuse high-resolution point spectral measurements with point clouds at close distances.
@@ II-C Robot Spectroscopy
The use of spectroscopy in robotics has gained traction as a way to classify household objects [25], survey geologic sites [26], explore underwater [27], and assess fruit ripeness [28]. Our previous work has been concerned with making pregrasp inferences using point spectral measurements [29, 30, 31]. However, in all these works we have assumed a fixed operating distance from object to sensor aperture, which simplified modeling of spectral signals. Further, the design and operation of these sensors assumed that a single-point scan was representative of the entire item composition, which is only true for select tasks in object manipulation.
To analyze spectral signatures of complex objects, we require precise spatial data. A robotic platform can position a point spectrometer at known orientations to construct a 3D map. For example, Perseverance’s X-Ray Fluorescence Spectrometer is positioned near a rock using a robotic arm, and a hexapod stage fine-tunes the sensor’s pose [32]. We seek to adapt this concept into an accessible platform for the general robotics community.
@@ III System Design
PROSPECT is a robot spectroscopy system that incorporates mechanical and optical elements for accurate formation and registration of spectral signatures with scanned geometries. PROSPECT uses a compact 6-degree-of-freedom (DoF) Stewart platform for precision orientation of a spectroscopy sensor suite relative to a measured surface. In this study, a 6-DoF robotic arm (UR3e, Universal Robotics) provides global motion around an object, while PROSPECT provides fine surface normalization and offsetting.
@@ III-A Mechanical
PROSPECT is designed to function as a precision end-effector adaptable to global motion platforms such as robot arms, gantries, or drones. PROSPECT contains a sensor suite with a fiber-optic cable, illuminating lamps, and time-of-flight sensors. These are mounted on the top platform of the Stewart mechanism, which allows for both spherical motion (for surface normal matching) and linear motion (for surface offsetting). This enables fine adjustment of the fiber-optic cable without the need to move the full weight of the global motion stage. The smoothness of spherical rotation also minimizes damage to the fiber optic cable that cannot be pinched or severely curved.
Figure 2: (Left) PROSPECT end-effector component diagram showing subcomponents. Individual motors and linkages are numbered. (Right) Identification of parameters and axes for kinematics and motion planning.
The PROSPECT end-effector is composed of a rigid base, a series of parallel links each composed of a horn $l_{h}$ and an arm $l_{a}$, leading to a sensorized platform as shown in Fig. 2. The base is 3D printed in Onyx on a Markforged X7 and contains wire routing and mounting to connect to a standard interface (Universal Robotics). Six servo motors (XL330-M288-T, Dynamixel) are mounted to the base. Each servo is oriented radially outward in a triangular pattern to maximize the motor packing density. The small angle between neighboring servos, $\theta_{b}$, is measured from the $z$ axis to the virtual point at the intersection between the servo axis and the plane of the arms.
The base also holds a small 3D printed electronics enclosure with a microcontroller (QtPy, Adafruit), a servo controller, and servo power board. A custom servo horn leading to the anchor point of the ball-ended arm is attached to each motor. Each ball end arm consists of one left and one right threaded rod end ball joint joined by an aluminum turnbuckle to allow adjustment of the lengths of the arms for different desired workspaces. The upper ball joint of each arm connects to the outer edge of the top platform, as shown in Fig. 2 below. The angle between the platform connection points, $\theta_{p}$, is measured from the z axis to the center of the upper ball joint.
@@ III-B Optical
The top platform contains the sensors and electronics that enable point spectroscopy. Current spectrometers are too fragile and bulky to be fully encapsulated in the top platform; instead, the center of the platform contains a shielded fiber optic cable (ThorLabs) to guide light from the field of view of the platform and into a spectrometer at the base of the manipulator. The cable contains a 400 $\mu$m silica glass core with a numerical aperture (NA) of 0.50. The top platform clamps the ferrule end of the fiber to a fixed normal orientation relative to the platform surface. The fiber optic cable from the top plate connects to a VNIR diffraction grating spectrometer (Pebble VIS-NIR, Ibsen Photonics) with sensitivity from 500-1100 nm in 256 discrete wavelength samples. PROSPECT pairs with any grating spectrometer operating in the VIS-SWIR range (400-2200 nm); thus, it is reconfigurable for sensitivity to either a broad or narrow range of the electromagnetic spectrum given the dictates of the inspection application.
Before operation, the spectrometer system is calibrated with a Spectralon reference standard and a dark noise collection according to [33].
@@
In (1), the tilde operator refers to the median of the spectral samples, collected over 10 measurements. The subscript $t$ indicates all measurements were collected with the same integration period. ${\tilde{L}}_{spec,t}$ refers to the median measurement taken from the Spectralon target at the nominal viewpoint distance. ${\tilde{D}}_{spec,t}$ refers to the median measurement taken with the spectrometer capped to incoming light. This reading accounts for dark current and other thermally induced electrical noise. ${S_{spec,t}}$ refers to the raw spectral signal for a particular sample and ${S_{calspec,t}}$ is the calibrated output.
The top platform contains 3 miniaturized Quartz Tungsten Halogen lamps (Ocean Insight). The lamps bathe the target surface in full-spectrum illumination, providing a sufficient signal for spectroscopic measurements. The lights operate at 5V and are controlled with a MOSFET circuit.
The top platform also contains 3 time-of-flight (ToF) distance sensors (STMicroelectronics VL53L4CD) oriented at 120 degree increments offset from the QTH lamps. The ToF sensors emit a pulse with a half-angle of 9. Given their spacing, the platform may use the three sensors to approximate a surface normal to the target surface at distances less than 6.0 cm. In this current work, these sensors are used to confirm the distance from the fiber ferrule tip to the object of interest.
@@ IV-A Inverse Kinematics
    Compute $\vec{q_{i}}$ from (3) and $\vec{l_{i}}$ from (4)
@@
    Solve for $\alpha_{i}$ using (5)
@@ IV-B Validation of Inverse Kinematics
To fully quantify the end-effector pose error, we report errors in position and orientation. Position error is given as the root mean squared error (RMSE) between the motion capture system’s observed 3D coordinate $(x_{1},y_{1},z_{1})_{k}$ and the commanded pose $(x_{2},y_{2},z_{2})_{k}$ for sample $k$ out of $n$ total randomly sampled poses. The RMSE metric is given as:
@@
Fig. 3 shows the results of the kinematics validation experiments. Overall, the system is found to have a kinematic accuracy of $3.73\pm 3.10$ mm in positional errors and $0.036\pm 0.025$ radians in rotational accuracy. This error arises primarily from motor and bearing backlash. The backlash’s effect on the platform varies based on its pose, but the positional and rotational error places an upper bound on the magnitude of the backlash of the system. The system’s minimal angular error indicates precise surface normal matching is possible.
@@ V Viewpoint Planning
With a point spectrometer, only a small portion of the object of interest is characterized per measurement. To characterize the entire object, we determine a subset of points that covers the surface of the object. We assume the existence of a point cloud that forms the basis for our viewpoint planning. To maximize scan consistency and spectral signal strength, the fiber optic cable is placed at a fixed distance from points of interest, as well as normal to the surface at that point.
@@ V-A Point Cloud Preprocessing
The raw point cloud from the 3D ToF Sensor (Azure Kinect, Microsoft) mounted above the arm contains points belonging the object of interest, as well as other extraneous points. Open3D [36] is used for the point cloud processing. The extraneous points are removed through a series of preprocessing steps. First, the cloud is cropped based off of empirically determined global bounds. Then, the table upon which the object rests is discarded by removing inliers belonging to a plane found with RANSAC [37]. Finally, noisy points are removed, and points belonging to the largest object remaining in the scene are selected using Density Based Clustering [38].
@@ V-B Approach Positioning
With the processed point cloud, viewpoints are selected using Algorithm 1. Consistently spaced points that cover the viewable surface are found through voxel-based uniform down-sampling. The normal is then estimated at each downsampled point. For each point and corresponding normal, the Stewart platform is positioned at a fixed distance from the surface along the normal such that the fiber optic probe is pointed parallel to the normal, facing the point. The arm is positioned at a point with only a Z-axis offset to the Stewart platform, with a fixed rotation for all points, so that Algorithm 1 generalizes to the possible configuration of a Stewart platform affixed to the end of a 3-DoF system such as an $XYZ$ Gantry. Our experiments use the MoveIt motion planning backend to achieve desired approach positions [39].